OpenStudy (anonymous):

An illumination flare is dropped over the edge of a 150 meter cliff. How long will it take for it to free fall to the bottom? Round your answer to the nearest tenth.

OpenStudy (jack1):

initial velocity = 0 distance = 150 acceleration = 9.8 m/s/s time = ???

OpenStudy (jack1):

s = ut + 1/2 at^2

OpenStudy (jack1):

initial velocity = 0 (u) distance = 150 (s) acceleration = 9.8 m/s/s (a) time = ??? (t)

OpenStudy (anonymous):

okay i think i got it